Methyl formate is a compound that is made up of 40.0% carbon, 6.70 % hydrogen, and 53.3 % oxygen by mass. Its molar mass is 60.05 g/mol. From this information, determine the empirical and molecular formulas of methyl formate.
(a) First of all, if you are given “percent by mass” data, and you know that “percent” means “parts per hundred”, what might you assume?
(b) Based on your assumption in part (a), calculate the lowest-whole-number ratio of moles of C, H, and O in the molecule (this is the “empirical” formula).
(c) What is the empirical mass (the molar mass of the empirical formula in part (b))?
(d) What is the ratio of molecular mass to empirical mass?
(e) What is the molecular formula, then?
Answer #1

a) First of all ,We will assume a fixed mass say M = 100g.

b) Mass of Carbon = (40*M)/100 = 40g

Mass of Hydrogen = (67*M)/100 = 6.7g

Mass of Oxygen = (53.3*M)/100 = 53.3g

Moles of C = 40/12 = 3.33 moles

Moles of H = 6.7/1 = 6.7 moles

Moles of O = 53.3/16 = 3.33 moles

Lowest ratio of C:H:O = 3.33:6.7:3.33 = 1:2:1

So, empirical formula = CH2O

c) Empirical mass = 12+2*1+16 = 30g/mol

d) Ratio of Molecular mass to empirical mass = 60.05/30 = 2

e) Molecular Formula = C2H4O2 or CHOOCH3